OpenStudy (anonymous):

3x+2y=-12 solve for y.

jimthompson5910 (jim_thompson5910):

First subtract 3x from both sides to get 2y=-3x-12 what is your next step?

OpenStudy (anonymous):

subtract 3x from both sides then divide to leave the y alone

OpenStudy (anonymous):

3x+2y=12 2y=-3x+12 y=-3x+12/2 y=-3x/2+6

OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12 -2y - 4 2y - 4 2y + 4 -2y + 4 help please